# A Frustratingly Simple Yet Highly Effective Attack Baseline: Over 90% Success Rate Against the Strong Black-box Models of GPT-4.5/4o/o1

### v1

@@ Abstract
Despite promising performance on open-source large vision-language models (LVLMs), transfer-based targeted attacks often fail against black-box commercial LVLMs. Analyzing failed adversarial perturbations reveals that the learned perturbations typically originate from a uniform distribution and lack clear semantic details, resulting in unintended responses. This critical absence of semantic information leads commercial LVLMs to either ignore the perturbation entirely or misinterpret its embedded semantics, thereby causing the attack to fail. To overcome these issues, we notice that identifying core semantic objects is a key objective for models trained with various datasets and methodologies. This insight motivates our approach that refines semantic clarity by encoding explicit semantic details within local regions, thus ensuring interoperability and capturing finer-grained features, and by concentrating modifications on semantically rich areas rather than applying them uniformly. To achieve this, we propose a simple yet highly effective solution: at each optimization step, the adversarial image is cropped randomly by a controlled aspect ratio and scale, resized, and then aligned with the target image in the embedding space. Experimental results confirm our hypothesis. Our adversarial examples crafted with local-aggregated perturbations focused on crucial regions exhibit surprisingly good transferability to commercial LVLMs, including GPT-4.5, GPT-4o, Gemini-2.0-flash, Claude-3.5-sonnet, Claude-3.7-sonnet, and even reasoning models like o1, Claude-3.7-thinking and Gemini-2.0-flash-thinking. Our approach achieves success rates exceeding 90% on GPT-4.5, 4o, and o1, significantly outperforming all prior state-of-the-art attack methods. Our optimized adversarial examples under different configurations are available at HuggingFace and training code at GitHub.
Figure 1: Example responses from commercial LVLMs to targeted attacks generated by our method.
@@ 1 Introduction
Adversarial attacks have consistently threatened the robustness of AI systems, particularly within the domain of large vision-language models (LVLMs) [26, 4, 41]. These models have demonstrated impressive capabilities on visual and linguistic understanding integrated tasks such as image captioning [36], visual question answering [30, 33] and visual complex reasoning [25, 34]. In addition to the progress seen in open-source solutions, advanced black-box commercial multimodal models like GPT-4o [1], Claude-3.5 [3], and Gemini-2.0 [37] are now extensively utilized. Their widespread adoption, however, introduces critical security challenges, as malicious actors may exploit these platforms to disseminate misinformation or produce harmful outputs. Addressing these drawbacks necessitates thorough adversarial testing in black-box environments, where attackers operate with limited insight into the internal configurations and training data of the models.
Current transfer-based approaches [43, 11, 15] typically generate adversarial perturbations that lack semantic structure, often stemming from uniform noise distributions with low success attacking rates on the robust black-box LVLMs. These perturbations fail to capture the nuanced semantic details that many LVLMs rely on for accurate interpretation. As a result, the adversarial modifications either go unnoticed by commercial LVLMs or, worse, are misinterpreted, leading to unintended and ineffective outcomes. This inherent limitation has motivated a deeper investigation into the nature and distribution of adversarial perturbations.
Our analysis reveals that a critical drawback in conventional adversarial strategies is the absence of clear semantic information within the perturbations. Without meaningful semantic cues, the modifications fail to influence the model’s decision-making process effectively. This observation is particularly relevant for commercial LVLMs, which have been optimized to extract and leverage semantic details from both local and global image representations. The uniform nature of traditional perturbations thus represents a significant barrier to achieving high attack success rates.
@@ 2.1 Large vision-language models
Transformer-based LVLMs integrate visual and textual modalities by learning rich visual-semantic representations from large-scale image-text datasets. These architectures power image captioning [36, 16, 8, 38], visual question answering [30, 33], and cross-modal reasoning [40, 31, 39]. Open-source models like BLIP-2 [23], Flamingo [2], and LLaVA [27] demonstrate strong generalization across vision-language benchmarks, while commercial models like GPT-4o, Claude-3.5 [3], and Gemini-2.0 [37], advance multimodal understanding with superior reasoning and real-world adaptation. Their black-box nature obscures vulnerability to adversarial perturbations, demanding systematic investigation of attack susceptibility.
2.2 LVLM Transfer-based adversarial attacks
Black-box adversarial attacks comprise query-based [10, 18] and transfer-based [9, 28] methods. Query-based approaches estimate gradients through multiple model interactions, incurring high computational costs. Transfer-based attacks generate adversarial samples on white-box surrogate models, exploiting cross-architecture transferability without querying the model. Commercial LVLMs challenge transfer-based attacks with unknown architectures, training data, and fine-tuned tasks, creating semantic gaps between surrogate and target models.
This paper investigates image-based adversarial attacks on LVLMs. AttackVLM [43] pioneered transfer-based targeted attacks using CLIP [35] and BLIP [23] as surrogates, demonstrating that image-to-image feature matching outperforms cross-modality optimization, a principle adopted by subsequent research [42, 15, 11]. CWA [7] and its successor SSA-CWA [11] extended this approach to examine commercial LVLMs robustness on Google’s Bard [37]. CWA enhances transferability through sharpness aware minimization [13, 6], optimizing ensemble models’ local optima flatness. SSA-CWA further incorporates spectrum transformation from SSA [29], achieving 22% success and 5% rejection rates on Bard.
AnyAttack [42] and AdvDiffVLM [15] leverage target image feature matching differently. AnyAttack uses it through large-scale self-supervised pre-training and dataset-specific fine-tuning. Despite some success on some commercial LVLMs, it produces template-like images with poor visual quality. AdvDiffVLM integrates feature matching into diffusion processes as guidance and implements Adaptive Ensemble Gradient Estimation (AEGE) for smoother ensemble scores. Its GradCAM-guided Mask Generation (GCMG) places perturbations in visually unimportant areas, enhancing imperceptibility. However, AdvDiffVLM achieves limited success against commercial LVLMs.
Figure 3: Heatmap visualization of perturbation aggregation using different crop schemes after various optimization steps. The scales control the range of proportions to the original image area.
3 Insights over Failed Attacks
Our investigation into failed adversarial transfers to target LVLMs in prior state-of-the-art solutions [43, 11, 42] reveals two key insights:
Uniform-like perturbation distribution. Analysis of failed adversarial samples in prior methods shows their perturbations follow a nearly uniform distribution, as in Fig. 5 and Fig. 3 first row. This uniform pattern indicates a lack of structural emphasis and a loss of diversity and details.
Vague description of black-box model. Next, we examine the outputs of various black-box VLMs on adversarial samples. We found that around 20% of the responses contained vague or ambiguous descriptions (e.g., “blurry”, “abstract”), as shown in Tab. 1. This indicates that, while the black-box model does detect something unusual in the image, it struggles to interpret it consistently and clearly.
@@ 3 Insights over Failed Attacks
AttackVLM [43] 6% 11% 45%
AnyAttack [42] 13% 13% 76%
SSA-CWA [11] 21% 29% 75%
@@
Details matter. The findings reveal that effective transferable adversarial samples require both semantic alignment and fine-grained visual details to successfully mislead target models. While semantic alignment provides the foundation, it is the subtle local features that carry the specific information needed to trigger misclassification. Current approaches that rely primarily on global similarity maximization struggle to preserve these crucial fine-grained details, limiting their effectiveness in generating transferable attacks.
@@ 4 Approach
[…]
Figure 4: Comparison of global similarity and ASR across different matching methods, including Global to Global, Local to Global and Local to Local.
Figure 5: Empirical cumulative distribution function of failed adversarial samples vs. uniform distribution. Shading shows standard deviation.
Framework Overview. Our approach aims to enhance the semantic richness within the perturbation by extracting details matching certain semantics in the target image. By doing so, we improve the transferability of adversarial examples across various models through a many-to-many/one matching, enabling them to remain effective against even the most robust black-box systems like GPT-4o, Gemini, and Claude. As shown in Fig. 2, at iteration $i$, the generated adversarial sample performs random cropping followed by resizing to its original dimensions. The cosine similarity between the local source image embedding and the global or local target image embedding is then computed using an ensemble of surrogate white-box models to guide perturbation updates. Through this iterative local-global or local-local matching, the central perturbed regions on the source image become progressively more refined, enhancing both semantic consistency and attack effectiveness, which we observe is surprisingly effective for commercial black-box LVLMs.
@@
Balancing Semantics and Consistency Between Feature and Image Spaces. Our local perturbation aggregation applied to the source image helps prevent an over-reliance on the target image’s semantic cues in the feature space. This is critical because the loss is computed directly from the feature space, which is inherently less expressive and does not adequately capture the intricacies of the image space. As shown in Fig. 5, we compare the global similarity between source and target images optimized using local and global perturbations. The Global-to-Global method achieves the highest similarity, indicating the best-optimized distance between the source and target. However, it results in the lowest ASR (i.e., worst transferability) on LVLMs, suggesting that optimized distance alone is not the key factor and that local perturbations on source can help prevent overfitting and enhance transferability.
By encoding enhanced semantic details through multiple overlapping steps, our method gradually builds a richer representation of the input. Meanwhile, the maintained consistency of these local semantic representations prevents them from converging into a uniform or homogenized expression. The combination of these enhanced semantic cues and diverse local expressions significantly improves the transferability of adversarial samples. Thus, we emphasize two critical properties for $\hat{\bf x}_{i}\in\mathcal{T}(\bf X)$:
@@
$\forall i,j,\quad\lvert\hat{\bf x}_{i}\cup\hat{\bf x}_{j}\rvert\geq\lvert\hat{\bf x}_{i}\rvert\ \text{and}\ \lvert\hat{\bf x}_{i}\cup\hat{\bf x}_{j}\rvert\geq\lvert\hat{\bf x}_{j}\rvert$ (4)
@@
Model Ensemble for Shared, High-quality Semantics. While our matching extracts detailed semantics, commercial black-box models operate on proprietary datasets with undisclosed fine-tuning objectives. Improving transferability requires better semantic alignment with these target models. We hypothesize that VLMs share certain semantics that transfer more readily to unknown models, and thus employ a model ensemble $\phi=\{f_{\phi_{1}},f_{\phi_{2}},\dots f_{\phi_{m}}\}$ to capture these shared elements. This approach formulates as:
@@
Our ensemble serves dual purposes. At a higher level, it extracts shared semantics that transfer more effectively to target black-box models. At a lower level, it can combine models with complementary perception fields to enhance perturbation quality. Models with smaller perception fields (e.g., transformers with smaller patch sizes) extract perturbations with finer details, while those with larger perception fields preserve better overall structure and pattern. This complementary integration significantly improves the final perturbation quality, as demonstrated in Fig. 6.
Training. To maximize $\mathcal{M}_{\mathcal{T}_{s},\mathcal{T}_{t}}$ while maintaining imperceptibility constraints, various adversarial optimization frameworks such as I-FGSM [22], PGD [32], and C&W [5], are applicable. For simplicity, we present a practical implementation that uses a uniformly weighted ensemble with I-FGSM, as illustrated in Algorithm 1. More formal and detailed formulations of the problem, along with derivations and additional algorithms, are provided in the Appendix.
Figure 6: 1) Left: visualization of perturbations generated by models with local-to-global matching. Numbers after ‘/’ indicate patch size. Models with smaller reception fields (14, 16) capture fine details, while larger ones (32) preserve better overall structure. The ensemble integrates these complementary strengths for high-quality perturbation. 2) Right: visualization of perturbation generated by other competitive methods. These perturbations are plotted with 5×5$\times$5 × magnitude, 1.5×1.5$\times$1.5 × sharpness and saturation for better visual effect.
@@ 5.1 Setup
We provide the experimental settings and competitive baselines below, with more details in the Appendix.
Victim black-box models and datasets. We evaluate three leading commercial multimodal large models: GPT-4.5, GPT-4o, o1, Claude-3.5-sonnet, Claude-3.7-sonnet, and Gemini-2.0-flash/thinking [37]. We use the NIPS 2017 Adversarial Attacks and Defenses Competition [20] dataset. Following [12], we sample 100 images and resize them to $224\times 224$ pixels. For enhanced statistical reliability, we then conduct evaluations on 1K images for the comparison with competitive methods in Sec. 5.3 in the Appendix.
@@
AdvDiffVLM [15] Ensemble 0.02 0.00 0.00 0.02 0.01 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.064 0.095
SSA-CWA [11] Ensemble 0.11 0.06 0.00 0.09 0.05 0.02 0.00 0.04 0.07 0.03 0.00 0.05 0.059 0.060
AnyAttack [42] Ensemble 0.44 0.20 0.04 0.42 0.46 0.21 0.05 0.48 0.25 0.13 0.01 0.23 0.048 0.052
@@
4 AttackVLM [43] 0.08 0.04 0.00 0.02 0.09 0.02 0.00 0.00 0.06 0.03 0.00 0.00 0.010 0.011
SSA-CWA [11] 0.05 0.03 0.00 0.03 0.04 0.03 0.00 0.04 0.03 0.02 0.00 0.01 0.015 0.015
AnyAttack [42] 0.07 0.02 0.00 0.05 0.10 0.04 0.00 0.05 0.03 0.02 0.00 0.02 0.014 0.015
@@
8 AttackVLM [43] 0.08 0.02 0.00 0.01 0.08 0.03 0.00 0.02 0.05 0.02 0.00 0.00 0.020 0.022
SSA-CWA [11] 0.06 0.02 0.00 0.04 0.06 0.02 0.00 0.06 0.04 0.02 0.00 0.01 0.030 0.030
AnyAttack [42] 0.17 0.06 0.00 0.13 0.20 0.08 0.01 0.14 0.07 0.03 0.00 0.06 0.028 0.029
@@
16 AttackVLM [43] 0.08 0.02 0.00 0.02 0.06 0.02 0.00 0.00 0.04 0.01 0.00 0.00 0.036 0.041
SSA-CWA [11] 0.11 0.06 0.00 0.09 0.05 0.02 0.00 0.04 0.07 0.03 0.00 0.05 0.059 0.060
AnyAttack [42] 0.44 0.20 0.04 0.42 0.46 0.21 0.05 0.48 0.25 0.13 0.01 0.23 0.048 0.052
@@
Surrogate models. We employ three CLIP variants [17] as surrogate models: ViT-B/16, ViT-B/32, and ViT-g-14-laion2B-s12B-b42K, for different architectures, training datasets, and feature extraction capabilities. We also include results on BLIP-2 [24] in the Appendix. Single-model method [43], if not specified, uses ViT-B/32 as its surrogate model. The ensemble-based methods [15, 42, 11] use the models specified in their papers.
Baselines. We compare against four recent targeted and transfer-based black-box attackers: AttackVLM [43], SSA-CWA [11], AnyAttack [42], and AdvDiffVLM [15].

### v2

@@ Abstract
Despite promising performance on open-source large vision-language models (LVLMs), transfer-based targeted attacks often fail against closed-source commercial LVLMs. Analyzing failed adversarial perturbations reveals that the learned perturbations typically originate from a uniform distribution and lack clear semantic details, resulting in unintended responses. This critical absence of semantic information leads commercial black-box LVLMs to either ignore the perturbation entirely or misinterpret its embedded semantics, thereby causing the attack to fail. To overcome these issues, we propose to refine semantic clarity by encoding explicit semantic details within local regions, thus ensuring the capture of finer-grained features and inter-model transferability, and by concentrating modifications on semantically rich areas rather than applying them uniformly. To achieve this, we propose a simple yet highly effective baseline: at each optimization step, the adversarial image is cropped randomly by a controlled aspect ratio and scale, resized, and then aligned with the target image in the embedding space. While the naïve source-target matching method has been utilized before in the literature, we are the first to provide a tight analysis, which establishes a close connection between perturbation optimization and semantics. Experimental results confirm our hypothesis. Our adversarial examples crafted with local-aggregated perturbations focused on crucial regions exhibit surprisingly good transferability to commercial LVLMs, including GPT-4.5, GPT-4o, Gemini-2.0-flash, Claude-3.5/3.7-sonnet, and even reasoning models like o1, Claude-3.7-thinking and Gemini-2.0-flash-thinking. Our approach achieves success rates exceeding 90% on GPT-4.5, 4o, and o1, significantly outperforming all prior state-of-the-art attack methods with lower $\ell_{1}/\ell_{2}$ perturbations. Our optimized adversarial examples under different configurations are available at HuggingFace and our training code at GitHub.
Figure 1: Examples from closed-source LVLMs to targeted attacks generated by our method.
@@ 1 Introduction
Adversarial attacks have consistently threatened the robustness of AI systems, particularly within the domain of large vision-language models (LVLMs) liang2024survey ; caffagni2024revolution ; zhang2024mm . These models have demonstrated impressive capabilities on visual and linguistic understanding integrated tasks such as image captioning image_caption_1 , visual question answering q_and_a ; q_and_a_2 and visual complex reasoning li2024enhancing ; park2025generalizing . In addition to the progress seen in open-source solutions, advanced black-box commercial multimodal models like GPT-4o gpt4 , Claude-3.5 claude , and Gemini-2.0 gemini are now extensively utilized. Their widespread adoption, however, introduces critical security challenges, as malicious actors may exploit these platforms to disseminate misinformation or produce harmful outputs. Addressing these drawbacks necessitates thorough adversarial testing in black-box environments, where attackers operate with limited insight into the internal configurations and training data of the models.
Current transfer-based approaches attackvlm ; dong2023robust ; advdiffusionvlm typically generate adversarial perturbations that lack semantic structure, often stemming from uniform noise distributions with low success attacking rates on the robust black-box LVLMs. These perturbations fail to capture the nuanced semantic details that many LVLMs rely on for accurate interpretation. As a result, the adversarial modifications either go unnoticed by commercial LVLMs or, worse, are misinterpreted, leading to unintended and ineffective outcomes. This inherent limitation has motivated a deeper investigation into the nature and distribution of adversarial perturbations.
Our analysis reveals that a critical drawback in conventional adversarial strategies is the absence of clear semantic information within the perturbations. Without meaningful semantic cues, the modifications fail to influence the model’s decision-making process effectively. This observation is particularly relevant for closed-source commercial LVLMs, which have been optimized to extract and leverage semantic details from both local and global image representations. The uniform nature of traditional perturbations thus represents a significant barrier to achieving high attack success rates.
@@ 2 Related Work
Large Vision-Language Models. Transformer-based LVLMs integrate visual and textual modalities by learning joint visual-semantic representations from large-scale image–text datasets. These models have underlaid core multimodal tasks such as image captioning image_caption_1 ; caption_2 ; caption_3 ; caption_4 , visual question answering q_and_a ; q_and_a_2 , and cross-modal reasoning corss_model_reason ; corss_model_reason1 ; corss_model_reason2 . Open-source LVLMs like BLIP-2 blip , Flamingo flamingo , and LLaVA llava demonstrate good capabilities on standard benchmarks, while closed-source systems such as GPT-4o gpt4 , Claude-3.7 claude , and Gemini-2.5 gemini exhibit better instruction-following, reasoning, and adaptation to real-world multimodal tasks. Despite these advances, the closed-source nature of commercial LVLMs conceals internal mechanisms and vulnerabilities, making it difficult to evaluate their robustness under adversarial scenarios. This calls for a systematic exploration of their susceptibility to carefully crafted input perturbations.
Transfer-Based Adversarial Attacks on LVLMs. Black-box attacks on LVLMs are either query-based blackbox_1 ; blackbox_2 , relying on repeated API access to estimate gradients, or transfer-based transfer_1 ; transfer_2 , which craft adversarial examples on surrogates without querying the target. While the latter is more efficient, transferability is hindered by the closed nature of commercial LVLMs, including undisclosed architectures and data, leading to significant semantic mismatches. Recent methods like AttackVLM attackvlm improve transfer success by aligning image-level features rather than cross-modal ones. This strategy influenced CWA cwa and SSA-CWA dong2023robust , which enhance transferability to models like Bard using sharpness-aware optimization and spectrum-based augmentation, achieving modest performance. Other approaches such as AnyAttack anyattack and AdvDiffVLM advdiffusionvlm explore self-supervised pretraining and diffusion-based generation, but still struggle against leading commercial LVLMs. These limitations highlight the need for more stable, semantically grounded gradient strategies, which our method aims to address.
Figure 3: Empirical cumulative distribution vs. uniform distribution on 20 randomly-sampled failed adversarial images. Shading shows standard deviation.
[…]
Figure 4: Comparison of global similarity and ASR across different matching schemes: Global to Global, Local to Global and Local to Local.
3 Investigations Over Failed Attacks
@@
AttackVLM attackvlm 6% 11% 45%
AnyAttack anyattack 13% 13% 76%
SSA-CWA dong2023robust 21% 29% 75%
@@
We investigate why prior state-of-the-art methods attackvlm ; dong2023robust ; anyattack have failed from two perspectives: 1) The perturbations from these methods tend to be uniformly distributed rather than highlighting statistically significant regions; 2) In many failed cases, the model does detect the perturbation but is unable to articulate detailed semantic content, resulting in vague or ambiguous descriptions. Some failed examples are provided in Appendix H.2.
Figure 5: Simulated heatmap visualization of perturbation aggregation across various steps using different crop schemes. The scales control the range of proportions to the original image area.
Uniform-like Perturbation Distribution. Fig. 4 and Fig. 5 (first row) illustrate that the perturbation in failed adversarial examples closely aligns with a uniform distribution, as indicated by the near-overlap between empirical cumulative distribution function (ECDF) and the ideal uniform CDF over 20 samples. The minimal deviation and tight standard deviation bands suggest that perturbations are spread evenly across the image space without preference for semantically meaningful regions. This uniform-like behavior implies a lack of targeted manipulation toward critical visual features, leading to weak semantic interference and ultimately ineffective attacks on LVLMs. In other words, the model perceives these perturbations as noise rather than meaningful semantic shifts.
Vague Description. To further validate that the model perceives these uniform perturbations as noise rather than meaningful semantic shifts, we quantify the proportion of vague descriptions. Specifically, we define vague descriptions as cases where the model uses terms like “blurry” or “abstract” to describe the detected artifacts or perturbations, instead of concrete semantic nouns. As shown in Tab. 1, while the black-box closed-source LVLMs do detect something unusual in the image, they struggle to interpret it consistently and clearly.
Similarity Trajectories. We further visualize the evolution of similarity trajectories during training to understand why local matching is less prone to overfitting compared to previous global matching strategies, and why it more effectively attacks LVLMs. As shown in Fig. 4, we observe that global representations lack sufficient randomness, causing the similarity (i.e., negative loss) to increase rapidly and saturate early. This early saturation limits further learning. In contrast, local matching converges more slowly, allowing the model to capture finer-grained details throughout training.
@@ 4 Approach
Framework Overview. Our approach aims to enhance the semantic richness within the perturbation by extracting details matching certain semantics in the target image. By doing so, we improve the transferability of adversarial examples through a many-to-many/one matching, enabling them to remain effective against even the most robust black-box systems like GPT-4o, Gemini, and Claude. As shown in Fig. 2, at iteration $i$, the generated adversarial sample performs random cropping followed by resizing to its original dimensions. The cosine similarity between the local source image embedding and the global or local target image embedding is then computed using an ensemble of surrogate white-box models to guide perturbation updates. The source-target pairs are randomly sampled. Through this iterative local-global or local-local matching, the central perturbed regions on the source image become progressively more refined, enhancing both semantic consistency and attack effectiveness, which we observe is surprisingly effective for commercial black-box LVLMs.
@@
Balancing Semantics and Consistency Between Feature and Image Spaces. Our local perturbation aggregation applied to the source image helps prevent an over-reliance on the target image’s semantic cues in the feature space. This is critical because the loss is computed directly from the feature space, which is inherently less expressive and does not adequately capture the intricacies of the image space. As shown in Fig. 4, we compare the global similarity between source and target images optimized using local and global perturbations. The Global-to-Global method achieves the highest similarity, indicating the best-optimized distance between the source and target. However, it results in the lowest ASR (i.e., worst transferability) on LVLMs, suggesting that optimized distance alone is not the key factor and that local perturbations on source can help prevent overfitting and enhance transferability. By encoding enhanced semantic details through multiple overlapping steps, our method gradually builds a richer representation of the input. Meanwhile, the maintained consistency of these local semantic representations prevents them from converging into a uniform or homogenized expression. The combination of these enhanced semantic cues and diverse local expressions significantly improves the transferability of adversarial samples. Thus, we emphasize two critical properties for $\hat{\bf x}_{i}\in\mathcal{T}(\bf X)$:
@@
$\forall i,j,\quad\lvert\hat{\bf x}_{i}\cup\hat{\bf x}_{j}\rvert>\lvert\hat{\bf x}_{i}\rvert\ \text{and}\ \lvert\hat{\bf x}_{i}\cup\hat{\bf x}_{j}\rvert>\lvert\hat{\bf x}_{j}\rvert$ (4)
@@
Model Ensemble for Shared, High-quality Semantics. While our matching extracts detailed semantics, commercial black-box models operate on proprietary datasets with undisclosed training objectives. Improving transferability requires better semantic alignment with these target models. We hypothesize that VLMs share certain semantics that transfer more readily to unknown models, and thus employ a model ensemble $\phi=\{f_{\phi_{1}},f_{\phi_{2}},\dots f_{\phi_{m}}\}$ to capture these shared elements. This approach formulates as:
@@
Our ensemble serves dual purposes. At a higher level, it extracts shared semantics that transfer more effectively to target black-box models. At a lower level, it can combine models with complementary perception sizes to enhance perturbation quality. Models with smaller receptive fields (e.g., transformers with smaller patch sizes) extract perturbations with finer details, while those with larger receptive fields preserve better overall structure and pattern. This complementary integration significantly improves the final perturbation quality, as demonstrated in Fig. 6.
Training. To maximize $\mathcal{M}_{\mathcal{T}_{s},\mathcal{T}_{t}}$ while maintaining imperceptibility constraints, various adversarial optimization frameworks such as I-FGSM ifgsm , PGD pgd , and C&W cw_attack , are applicable. For simplicity, we present a practical implementation that uses a uniformly weighted ensemble with I-FGSM, as illustrated in Algorithm 1. More formal and detailed formulations of the problem, along with derivations and additional algorithms, are provided in the Appendix.
@@ 5.1 Setup
We provide the experimental settings and strong baselines below, with more details in the Appendix.
Victim Black-box Models and Datasets. We evaluate three leading commercial black-box multimodal large language model families: GPT-4.5, GPT-4o, o1, Claude-3.5-sonnet, Claude-3.7-sonnet, and Gemini-2.0-flash/thinking gemini . We use the NIPS 2017 Adversarial Attacks and Defenses Competition nips-2017-defense-against-adversarial-attack dataset. Following dong2023robust , we sample 100 images and resize them to $224\times 224$ pixels. For enhanced statistical reliability, we then conduct evaluations on 1K images for the comparison with competitive methods in Sec. G.1 of the Appendix. Our source-target image training pairs are randomly sampled.
Surrogate Models. We employ three CLIP variants opanai_clip_software as surrogate models: ViT-B/16, ViT-B/32, and ViT-g-14-laion2B-s12B-b42K, for different architectures, training datasets, and feature extraction capabilities. We also include results on BLIP-2 blip2 in the Appendix. Single-model method attackvlm , if not specified, uses ViT-B/32 as its surrogate model. The ensemble-based methods advdiffusionvlm ; anyattack ; dong2023robust use the models specified in their papers.
Baselines. We compare against four recent targeted and transfer-based black-box attackers: AttackVLM attackvlm , SSA-CWA dong2023robust , AnyAttack anyattack , and AdvDiffVLM advdiffusionvlm .
Figure 6: 1) Left: visualization of perturbations generated by models with local-to-global matching. Numbers after ‘/’ indicate patch size. Models with smaller receptive fields (14, 16) capture fine details, while larger ones (32) preserve better overall structure. The ensemble integrates these complementary strengths for high-quality perturbation. 2) Right: visualization of perturbation generated by other competitive methods. These perturbations are plotted with $5\times$ magnitude, $1.5\times$ sharpness and saturation for better visual effect.
@@
AdvDiffVLM advdiffusionvlm Ensemble 0.02 0.00 0.00 0.02 0.01 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.064 0.095
SSA-CWA dong2023robust Ensemble 0.11 0.06 0.00 0.09 0.05 0.02 0.00 0.04 0.07 0.03 0.00 0.05 0.059 0.060
AnyAttack anyattack Ensemble 0.44 0.20 0.04 0.42 0.46 0.21 0.05 0.48 0.25 0.13 0.01 0.23 0.048 0.052
@@
4 AttackVLM attackvlm 0.08 0.04 0.00 0.02 0.09 0.02 0.00 0.00 0.06 0.03 0.00 0.00 0.010 0.011
SSA-CWA dong2023robust 0.05 0.03 0.00 0.03 0.04 0.03 0.00 0.04 0.03 0.02 0.00 0.01 0.015 0.015
AnyAttack anyattack 0.07 0.02 0.00 0.05 0.10 0.04 0.00 0.05 0.03 0.02 0.00 0.02 0.014 0.015
@@
8 AttackVLM attackvlm 0.08 0.02 0.00 0.01 0.08 0.03 0.00 0.02 0.05 0.02 0.00 0.00 0.020 0.022
SSA-CWA dong2023robust 0.06 0.02 0.00 0.04 0.06 0.02 0.00 0.06 0.04 0.02 0.00 0.01 0.030 0.030
AnyAttack anyattack 0.17 0.06 0.00 0.13 0.20 0.08 0.01 0.14 0.07 0.03 0.00 0.06 0.028 0.029
@@
16 AttackVLM attackvlm 0.08 0.02 0.00 0.02 0.06 0.02 0.00 0.00 0.04 0.01 0.00 0.00 0.036 0.041
SSA-CWA dong2023robust 0.11 0.06 0.00 0.09 0.05 0.02 0.00 0.04 0.07 0.03 0.00 0.05 0.059 0.060
AnyAttack anyattack 0.44 0.20 0.04 0.42 0.46 0.21 0.05 0.48 0.25 0.13 0.01 0.23 0.048 0.052